OpenStudy (anonymous):

Help with proving whether or not the equation is a trig identity? Medal and fan. I need help proving whether or not cos^2x(1+tan2x)=1 is an identity Thanks(:

OpenStudy (loser66):

yes, it is open the bracket, what do you have?

OpenStudy (anonymous):

1-cos(2x)/2 *(1 + 1-cos(2x)/1+cos(2x)

OpenStudy (loser66):

\[cos^2(x)(1+tan^2(x))= cos^2(x) + cos^2(x) tan^2(x)\] now replace \(tan ^2 (x) = \dfrac{sin^2(x)}{cos^2(x)}\) what do you have?

OpenStudy (anonymous):

1-cos(2x)/ 1+ cos(2x) right?

OpenStudy (loser66):

nope. \[cos^2(x)(1+tan^2(x))= cos^2(x) + cos^2(x) tan^2(x)=cos^2(x) +cos^2(x)\dfrac{sin^2(x)}{cos^2(x)}\\= cos^2(x)+sin^2(x)=1\] that is identity

OpenStudy (anonymous):

Oh, I was going to far. I was giving the identities of the sin^2(x) and cos^2(x) sorry. Thanks, I understand that(: